[Brahms-l] Shift report 20050514 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat May 14 2005 - 15:49:46 EDT
Shifter leader was Thomas Materna
Shifters were 


Beam ok (yellow intensity stable), beam dumped at ~12:00 and back at ~15:00
runs 14634 to 14642, mostly physics ON.
Ramiro thinks there is a gain drift in the drift chambers. After the beam was back, we made to short runs with different scaledown factors for tr2, just to check the possibility of a problem with the readout when the rate is high. Ramiro is working on it...

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat May 14 15:50:01 2005

This archive was generated by hypermail 2.1.8 : Sat May 14 2005 - 15:50:14 EDT